stlectric dipole made of 1 z and 1.2 uc held at ananometer apartis putin an electric field of How much torque needed to take this arrangement to go with respect to the directien of the b How much work is done as it rotates back e are e two in
Answer #1

Solution: Toque Sin d=10-9 m E-33N/C q=1.2×10-6 C t 1.2x10-6 x109 x33sin 90 3.91x104 Nm Work done θ = 90